Circular No. 5999

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER 4 (1994k) Ephemeris from the orbital elements on IAUC 5998:

1994TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
May 28 17^{h}14.^{m}10 −10^{o}47.^{′}8 2.152 3.135 163.^{o}0 5.^{o}4 16.6
June 7 17 06.99 −11 47.8 2.109 3.111 168.9  3.6 16.5
17 16 59.52 −12 57.4 2.093 3.087 165.7  4.7 16.5
27 16 52.48 −14 14.7 2.104 3.065 156.8  7.5 16.5
July 7 16 46.59 −15 37.3 2.142 3.044 146.9 10.5 16.5
17 16 42.45 −17 03.0 2.203 3.025 136.9 13.3 16.5
27 16 40.45 −18 29.5 2.283 3.008 127.2 15.6 16.6

4U 0115+634 R. B. Wilson, M. H. Finger, and D. M. Scott report for the Compton

Gamma Ray Observatory BATSE pulsar team: “The outburst from the xray binary pulsar 4U 0115+634 previously reported on IAUC 5990 continues as of May 30. The flux in the energy band 20-50 keV rose at a constant rate from the onset of detection on May 8 to ∼ 65 mCrab on May 18, increased at a lesser rate to ∼ 75 mCrab on May 28, then increased more rapidly, reaching ∼ 180 mCrab on May 30.”

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) D. Jewitt and N. Trentham, University of Hawaii, communicate: “New

observations from the University of Hawaii 2.2-m telescope show that nucleus 15 (= G) is double. On May 7 UT, the fainter component (15b) appears 3′′.5 east and 3′′.7 north of the brighter (15a). The red magnitude difference measured in a 1′′-radius aperture is 3.8. The fainter component is not evident in observations taken in previous months under similar observing conditions.”

SUPERNOVA 1994I IN NGC 5194 L. V. Morrison and R. W. Argyle, Royal Greenwich Observatory, report

the following V magnitudes, derived from the observations mentioned on IAUC 5989 (the accuracy ranges from ±0.10 mag at V = 13.0 to ±0.20 at 14.0): Apr. 11.06 UT, 12.93; 12.06, 12.90; 13.05, 13.12; 14.05, 13.10; 15.05, 13.27; 16.04, 13.33; 23.03, 13.98; 25.02, 13.88.

1994 June 1 Daniel W. E. Green